# Relaying Quantum Information

### v1

@@ IV-C Unlimited Entanglement Assistance
In the upcoming sections, we provide a detailed proof for the achievable rates in Theorem 2 and Theorem 3.
V Analytic Tools: The FQSW Theorem
Before presenting the achievability proof, we provide a quick overview of the fully quantum Slepian-Wolf (FQSW) protocol as presented in Ref. [49]. This summary broadly follows the one given in Ref. [48]. Consider a pure state ${|{\psi}\rangle}_{ABC}$ shared by Alice, Bob, and a reference system held by Charlie. Now, Alice wishes to send her state to Bob with high fidelity and use as few qubits as possible. In the FQSW protocol, Alice first applies a unitary $U_{A\rightarrow A_{1}A_{2}}$, drawn according to the Haar measure, and then sends $A_{2}$ to Bob.
The FQSW result is a decoupling theorem. The protocol induces a state such that the $A_{1}$ and $C$ sub-systems are “almost" in a product state at the end of the protocol. As shown in Ref. [49], one may evaluate how close the resulting state is to a decoupled state by:
$\mathbb{E}_{U}\left\|\rho_{A_{1}C}-\frac{\mathbbm{1}_{A_{1}}}{|A_{1}|}\otimes\psi_{C}\right\|^{2}_{1}\leq\frac{|A||C|}{|A_{2}|^{2}}\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]$ (11)
where the expectation is taken over the random unitary $U_{A\rightarrow A_{1}A_{2}}$, according to the Haar measure. Applying this protocol to a block of $n$ transmissions from Alice to Bob, we have $\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]\approx 2^{-nH(AC)_{\psi}}$ for sufficiently large $n$, by the Quantum AEP Property [50].
Since Eq. (11) holds on average, there must exist at least one unitary that satisfies the bound. That is, using the state-approximation notation, there exists $U_{A\rightarrow A_{1}A_{2}}$ such that
with the proximity parameter $\varepsilon=\frac{|A||C|}{|A_{2}|^{2}}\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]$.
Since Bob holds the purifying system of $A_{1}C$, there exists a local unitary $U^{\prime}$ that splits his purifying system into two separate purifying systems of the two sub-systems $A_{2}$ and $C$, respectively. The purifying system of $C$ is exactly the system that Alice wanted to send, $A_{2}$, together with Bob’s system $B$. Whereas, $A_{1}$ becomes half of a maximally entangled pair, shared by Alice and Bob. Therefore, as a by-product of the FQSW protocol, we obtain an entanglement resource shared by Alice and Bob at the end of the protocol.
VI Proof of Proposition 1
First, we introduce the notation for the entanglement resources consumed and generated during the protocol. The quantum systems $G_{E}$ and $G_{B}$ represent the pre-shared entanglement resources available at the relay and the destination receiver (Bob), respectively, with associated consumption rates defined as $L_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|G_{E}|$ and $L_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|G_{B}|$.
In accordance with Sec. V, the quantum systems $\hat{G}_{1,A}$ and $\hat{G}_{2,A}$ represent the by-product entanglement generated between the sender (Alice) and the relay, and between the sender and the destination receiver, respectively. These systems are located at the sender, with corresponding generation rates given by $\hat{L}_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|\hat{G}_{1,A}|$ and $\hat{L}_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|\hat{G}_{2,A}|$.
VI-A Decoupling Properties
We begin with decoupling properties. Let ${|{\sigma_{A_{1}AD}}\rangle}$.
VI-A1 Decoupling at the relay’s transmitter
Consider the product state, ${|{\sigma_{A_{1}AD}}\rangle}^{\otimes n}$, where $A_{1}$ is the encoder ancilla, $A$ the channel input, and $D$ the relay output. Projecting onto the typical subspace of $A_{1}$ gives:
${|{\psi_{A_{1,\text{typ}}^{n}A^{n}D^{n}}}\rangle}=(\Pi_{A_{1}^{n}\to A_{1,\text{typ}}^{n}}\otimes\mathbbm{1}_{AD}){|{\sigma_{A_{1}AD}}\rangle}^{\otimes n}.$ (13)
As in Sec. V, we apply a random unitary $U_{A_{1,\text{typ}}^{n}\to RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}}$, where $R$ is the reference system with $|R|=2^{nQ}$. The resulting state is
${|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}=\left(U\otimes\mathbbm{1}_{A^{n}D^{n}}\right){|{\psi_{A_{1,\text{typ}}^{n}A^{n}D^{n}}}\rangle}.$ (14)
Considering the reduced state $\theta_{RG_{E}G_{B}D^{n}}$, we can “decouple" between $D^{n}$ and $RG_{E}G_{B}$, and write:
$\theta_{RG_{E}G_{B}D^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\frac{\mathbbm{1}_{RG_{E}G_{B}}}{|RG_{E}G_{B}|}\otimes\sigma_{D^{n}},$ (15)
which, by the FQSW Theorem (see Sec. V), holds for
$\varepsilon_{1}\equiv\varepsilon_{1}(n)$ $=\frac{|A_{1,\text{typ}}^{n}|\cdot|D^{n}|}{|\hat{G}_{1,A}|^{2}|\hat{G}_{2,A}|^{2}}\cdot 2^{-nH(A_{1}D)}$
$\leq 2^{-n(H(A_{1}|D)_{\sigma}-L_{E}-L_{B}+\hat{L}_{E}+\hat{L}_{B}-Q-\delta)},$ (16)
$Q<H(A_{1}|D)_{\sigma}-L_{E}-L_{B}+\hat{L}_{E}+\hat{L}_{B}-\delta.$ (17)
VI-A2 Decoupling at the relay’s receiver
${|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle}=\\ (\mathbbm{1}\otimes\left(V^{\mathcal{N}}_{AD\to BEJ_{O}})^{\otimes n}\right){|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle},$ (18)
where $R$ is Alice’s resource, $E^{n}$ is received at the relay, $B^{n}$ at the destination receiver, and $J_{O}^{n}$ represents the environment of the joint output, $B^{n}E^{n}$. We denote the composite environment of the relay by $J_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}BJ_{O}$, and consider the reduced state $\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}J_{E}^{n}}$. Decoupling the relay’s environment:
$\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}J_{E}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\frac{\mathbbm{1}_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}}}{|R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}|}\otimes(\mathcal{N}_{A\to J_{E}}({|{\sigma_{A}}\rangle}))^{\otimes n},$ (19)
where
$\varepsilon_{2}(n)$ $=\frac{|A_{1,typ}||J_{E}^{n}|}{|G_{E}|^{2}}\cdot 2^{-nH(A_{1,typ}J_{E})}$
$\leq 2^{-n(I(A_{1}\rangle E)_{\omega}+L_{E}-L_{B}-\hat{L}_{E}-\hat{L}_{B}-Q-\delta)},$ (20)
for sufficiently large $n$, where the inequality follows from the same arguments as used in Sec. VI-A1. The last expression tends to zero as $n\to\infty$, if
$Q<I(A_{1}\rangle E)_{\omega}+L_{E}-L_{B}-\hat{L}_{E}-\hat{L}_{B},$ (21)
with respect to $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$.
VI-A3 Decoupling at the destination receiver
We now consider the overall transmission from Alice to Bob, treating the relay’s output as part of Bob’s environment. The output state is as in Eq. (18). Denote Bob’s environment by $J_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}EJ_{O}$. Similarly as before, decoupling Bob’s environment yields
$\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{E}J_{B}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{3}}}}{{\approx}}}\frac{\mathbbm{1}_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{E}}}{|R\hat{G}_{A}G_{E}|}\otimes(\mathcal{N}_{A\to J_{B}}({|{\sigma_{A}}\rangle}))^{\otimes n},$ (22)
where
$\varepsilon_{3}(n)$ $=\frac{|A_{1,\mathrm{typ}}||J_{B}^{n}|}{|G_{B}|^{2}}\cdot 2^{-nH(A_{1,\mathrm{typ}}J_{B})}$
$\leq 2^{-n(I(A_{1}\rangle B)_{\omega}-L_{E}+L_{B}-\hat{L}_{E}-\hat{L}_{B}-Q-\delta)},$ (23)
for sufficiently large $n$. This tends to zero if
$Q<I(A_{1}\rangle B)_{\omega}-L_{E}+L_{B}-\hat{L}_{E}-\hat{L}_{B}.$ (24)
VI-B Encoding and decoding procedures
Throughout, we use the following convention for block indexing to clearly distinguish between the current and past instances of a quantum system during the $T$-block encoding and decoding process. A quantum system labeled as $S$ refers to the system at block index $j$, i.e., $S\stackrel{{\scriptstyle\mathrm{def}}}{{=}}S(j)$. A quantum system labeled as $S^{\prime}$ refers to the system at the previous block, i.e., $S^{\prime}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}S(j-1)$.
VI-B1 Transmitter, encoder and channel isometries
For each block $j$, there exists a family of partial isometries $W_{MG_{A}\rightarrow A^{n}\hat{G}_{1,A}\hat{G}_{2,A}}$ such that, according to the Sec. VI-A1, the following approximation holds:
${|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\\ \left(W_{MG_{1,A}G_{2,A}\rightarrow A^{n}\hat{G}_{1,A}\hat{G}_{2,A}}\otimes\mathbbm{1}\right)\cdot\\ \left(\Phi_{RM}\otimes\Phi_{G_{1,A}G_{E}}\otimes\Phi_{G_{2,A}G_{B}}\otimes\sigma_{D^{n}}\right),$ (25)
where $\Phi_{RM}$ is the maximally entangled message state between the reference $R$ and the “message" register $M$, and $\Phi_{G_{1,A}G_{E}}\otimes\Phi_{G_{2,A}G_{B}}$ is the entanglement assistance shared between Alice and the relay, and between Alice and Bob. The output systems $A^{n}\hat{G}_{1,A}\hat{G}_{2,A}$ jointly purify the reduced state $\theta_{RG_{E}G_{B}D^{n}}$.
The channel output is then
${|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle}=\\ \left[(V^{\mathcal{N}}_{AD\to BEJ_{O}})^{\otimes n}\otimes\mathbbm{1}\right]{|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}.$ (26)
VI-B2 Relay encoder and decoder
We begin with the relay encoder. For each block $j-1$, there exists a family of partial isometries $\Lambda_{\tilde{M}^{\prime}\hat{G}_{E}^{\prime n}\rightarrow{D}^{n}G_{E}^{\prime}}$ such that the following approximation holds:
${|{\theta_{R^{\prime}G_{E}^{\prime}G_{B}^{\prime}\hat{G}_{1,A}^{\prime}\hat{G}_{2,A}^{\prime}A^{\prime n}{D}^{n}}}\rangle}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\left[\Lambda_{\tilde{M}^{\prime}\hat{G}_{E}^{\prime n}\rightarrow{D}^{n}G_{E}^{\prime}}\otimes\mathbbm{1}\right]\\ \cdot\Phi_{R^{\prime}\tilde{M}^{\prime}}\otimes\Phi_{\hat{G}_{1,A}^{\prime}\hat{G}_{E}^{\prime n}}\otimes\Phi_{\hat{G}_{2,A}^{\prime}\hat{G}_{B}^{\prime n}}\otimes\psi_{B^{\prime n}J_{O}^{\prime n}},$ (27)
where $\theta_{D^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\sigma_{D}^{\otimes n}$, based on Sec. VI-A2. This is important for the next stage.
As for the relay decoder, invoking the results of Sec. VI-A2, the partial isometries $\Gamma_{E^{n}G_{E}\rightarrow\tilde{M}\hat{G}_{E}^{n}}$ satisfy
$\Phi_{R\tilde{M}}\otimes\Phi_{\hat{G}_{1,A}\hat{G}_{E}^{n}}\otimes\psi_{\hat{G}_{2,A}G_{B}B^{n}J_{O}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\\ \left[\Gamma_{E^{n}G_{E}\rightarrow\tilde{M}\hat{G}_{E}^{n}}\otimes\mathbbm{1}\right]\cdot{|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle},$ (28)
where the quantum system $\tilde{M}$ is the quantum “message restoration" system of the relay.
VI-B3 Destination decoder
Similarly to the above, according to Sec. VI-A3, for each block $j$, there exists a family of partial isometries $\Delta_{B^{n}G_{B}\rightarrow\hat{M}\hat{G}_{B}^{n}}$ such that:
$\Phi_{R\hat{M}}\otimes\Phi_{\hat{G}_{2,A}\hat{G}_{B}^{n}}\otimes\psi_{\hat{G}_{1,A}G_{E}E^{n}J_{O}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{3}}}}{{\approx}}}\\ \left[\Delta_{B^{n}G_{B}\rightarrow\hat{M}\hat{G}_{B}^{n}}\otimes\mathbbm{1}\right]\cdot{|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle},$ (29)
where the quantum system $\hat{M}$ is the quantum “message estimation" system of the destination receiver, producing $\approx\Phi_{R\hat{M}}$ between Alice and Bob. ∎
VII Proof of Theorem 2 and Theorem 3
In this section we prove Theorem 2 and Theorem 3 based on our result in Proposition 1.
VII-A Unassisted case
Taking the entanglement assistance rates to be $L_{E}=\hat{L}_{E}\rightarrow 0$ and $L_{B}=\hat{L}_{B}\rightarrow 0$, and combining the constraints in Eq. (17), Eq. (21), and Eq. (24) yields the following bound:
The inequality in Eq. (17) is inactive, since $H(A_{1}|D)_{\sigma}=I(A_{1}\rangle A)_{\sigma}$ by entropy duality, and $I(A_{1}\rangle A)_{\sigma}\geq I(A_{1}\rangle B)_{\omega}$ by the data processing inequality. This concludes the proof of Theorem 2.
VII-B Unlimited assistance
Assuming that only Bob has access to entanglement assistance, we set the relay’s entanglement rates to be $L_{E}=\hat{L}_{E}\rightarrow 0$. Isolating $Q$ in Eq. (24) and Eq. (17), we achieve
$Q<\frac{1}{2}I(A_{1};B)_{\omega}-\frac{1}{2}I(A_{1};D)_{\sigma}-\delta.$ (31)
This concludes the proof. ∎
VIII Summary and Discussion
We study quantum information transmission over a three-terminal quantum relay network, focusing on the decode-forward coding strategy in both entanglement-assisted and unassisted settings. Our model captures key tradeoffs between repeater-assisted and repeaterless communication, and highlights the role of entanglement between the transmitter and the relay. Furthermore, our framework generalizes previous results, which focus on either repeaterless transmission or repeater-aided communication. These insights serve as a stepping stone towards the design of quantum networks (in the long-term, a quantum Internet), and provide a foundation for practical quantum repeaters and long-distance QKD.
One avenue for future research is to develop alternative coding strategies. In particular, we may require the relay to decode only a part of the information. This is particularly relevant when the direct channel from Alice to Bob is not strictly noisier (e.g., degraded) than the channel to the relay. Classical information theory suggests that there is often an optimal balance between (a) the amount of information decoded through the relay and (b) the amount decoded independently by the destination receiver [52, 35]. Exploring how this trade-off manifests in the quantum setting may yield new insights into quantum information transmission over relay channels.

### v2

@@ IV-D Entanglement-Assisted Capacity
In the upcoming sections, we provide interpretation and a concrete example to further illustrate the results for the achievable rates in Theorem 2, Theorem 3 and Theorem 4.
V Interpretation
In what follows we explain the intuition behind each of the three conditions given by Eq. (1) in Proposition 1.
V-1 Full Decode-Forward Interpretation
First, suppose that the relay decodes the entire message. In this case, we set $Q^{\prime\prime}\to 0$ and $L^{\prime\prime}_{B}=\hat{L}^{\prime\prime}_{B}\to 0$. We also set $A_{2}$ to be degenerate (say, a register of dimension $1$). Next, we define the net entanglement consumption of EPR pairs as $\Delta L^{\prime}_{B}=L^{\prime}_{B}-\hat{L}^{\prime}_{B}$ and $\Delta L^{\prime}_{E}=L^{\prime}_{E}-\hat{L}^{\prime}_{E}$, shared between Alice and Bob and between Alice and the relay, respectively. Under these assumptions, Eq. (1) yields
$\begin{split}&Q^{\prime}+\Delta L^{\prime}_{B}+\Delta L^{\prime}_{E}\leq H(A_{1}|D)_{\sigma}\\ &Q^{\prime}-\Delta L^{\prime}_{E}\leq I(A_{1}\rangle E)_{\omega}-(L^{\prime}_{B}+\hat{L}^{\prime}_{B})\\ &Q^{\prime}-\Delta L^{\prime}_{B}\leq I(A_{1}\rangle B)_{\omega}-(L^{\prime}_{E}+\hat{L}^{\prime}_{E}).\end{split}$ (15)
Then, the first condition in Eq. (15) states that the quantum information transmitted by Alice, together with the entanglement consumed from her share of the EPR pairs, must fit within the channel input available to Alice, conditioned on the relay’s input.
In a Block Markov coding scheme, Alice transmits a sequence of blocks $A^{n}(1),\ldots,A^{n}(T)$, each of size $n$, in order to distribute a “quantum message" sequence $M(2),\ldots,M(T)$, each of dimension $2^{nQ}$. The transmission in each block not only carries new information but also forwards part of the information from the previous block.
The second condition in Eq. (15) bounds the net information rate $Q^{\prime}-\Delta L_{E}^{\prime}$, having accounted for the entanglement resources with the relay. The subtraction of $(L^{\prime}_{B}+\hat{L}^{\prime}_{B})$ reflects the bottleneck introduced by the relay. This arises due to the nature of our Block Markov coding scheme, where the entanglement assistance available to Bob in block $j$ is effectively generated by the relay in block j1j-1.
The third condition shown in Eq. (15) has a similar structure. Here, we subtract $(L^{\prime}_{E}+\hat{L}^{\prime}_{E})$, again reflecting the relay bottleneck. In addition, $\Delta L^{\prime}_{B}$ has a negative sign (opposite to the first condition), corresponding to the net entanglement consumption between Alice and Bob.
V-2 Partial Decode-Forward Interpretation
The above discussion generalizes to the case where part of the information is sent without the relay’s help, i.e., $Q^{\prime\prime}>0$. Recall that our coding scheme includes two message registers $M^{\prime}$ and $M^{\prime\prime}$, of rates $Q^{\prime}$ and $Q^{\prime\prime}$, respectively, as presented in Eq. (1). The relay decodes $M^{\prime}$ alone, and does not decode $M^{\prime\prime}$. In this case, we define $\Delta L^{\prime\prime}_{B}=L^{\prime\prime}_{B}-\hat{L}^{\prime\prime}_{B}$ as the net entanglement consumption in the encoding of $M^{\prime\prime}$.
Based on the intuition given above for the Eq. (15), the Eq. (15) changes to
$\begin{split}&Q^{\prime}+\Delta L^{\prime}_{B}+\Delta L^{\prime}_{E}+\Delta L^{\prime\prime}_{B}\leq H(A_{1}|D)_{\sigma}\\ &Q^{\prime}-\Delta L^{\prime}_{E}\leq I(A_{1}\rangle E)_{\omega}-(L^{\prime}_{B}+\hat{L}^{\prime}_{B})\\ &Q^{\prime}-\Delta L^{\prime}_{B}-\Delta L^{\prime\prime}_{B}\leq I(A_{1}\rangle B)_{\omega}-(L^{\prime}_{E}+\hat{L}^{\prime}_{E}).\end{split}$ (16)
where only the first and third conditions now include the additional term $\Delta L^{\prime\prime}_{B}$, accounting for the entanglement consumed for the encoding of $M^{\prime\prime}$, and the net entanglement consumption in the decoding of $M^{\prime\prime}$, respectively.
VI Example: Quantum erasure channel
Figure 4: A quantum relay channel with orthogonal components considered in our example. Alice’s (sender’s) input AA splits into $A^{\prime}$ and $A^{\prime\prime}$, entangled with auxiliaries $A_{1}$ and $_{2}. AA^{\prime}$ is sent through a relay-assisted link, while $A^{\prime\prime}$ is sent directly to Bob (receiver). The receiver obtains the outputs $B^{\prime}$ and $B^{\prime\prime}$, respectively.
To illustrate our results, we consider the quantum erasure channel, a widely used model in quantum communication [65, 66]. An erasure channel with erasure probability $p$ either transmits the input quantum state perfectly with probability 1p1-p, or replaces it with an “erasure” state $|{e}\rangle$} with probability $p$. Formally, for an input state $\rho$ on a qudit space $\mathbb{C}^{d}$ of dimension $d$, the channel acts as
$\mathcal{E}^{(p)}(\rho)=(1-p)\,\rho+p\,{|{e}\rangle}\!{\langle{e}|},$
where $|{e}\rangle$} is orthogonal to the qudit space $\mathbb{C}^{d}$. Hence, the output dimension is $d+1$.
We consider a quantum relay channel that consists of orthogonal components, as shown in Fig. 4. Specifically, Alice’s input $A$ consists of two components, $A^{\prime}$ and $A^{\prime\prime}$, which are sent through separate links. The first link is repeater aided and goes through the relay, whereas the latter is a direct repeaterless link to Bob. The system at the destination receiver, Bob, consists of the respective outputs, $B^{\prime}$ and $B^{\prime\prime}$. We assume that $A^{\prime},A^{\prime\prime},D,B^{\prime},B^{\prime\prime},E$ are all single qubit registers. The overall quantum channel is
$\mathcal{N}_{A^{\prime}A^{\prime\prime}D\to B^{\prime}B^{\prime\prime}E}=(\mathcal{E}^{(\beta)}_{B^{\prime}\to E}\circ\mathcal{E}^{(\alpha)}_{A^{\prime}D\to B^{\prime}})\otimes\mathcal{E}^{(\gamma)}_{A^{\prime\prime}\to B^{\prime\prime}}$. (17)
Naturally, we choose the ancilla $A_{1}$ to be maximally entangled with the relay-link input, $A^{\prime}$, and the ancilla $A_{2}$ to be maximally entangled with the direct-link input, $A^{\prime\prime}$. The initial state is thus taken as ${|{\sigma}\rangle}={|{\Phi_{A_{1}A^{\prime}D}}\rangle}\otimes{|{\Phi_{A_{2}A^{\prime\prime}}}\rangle}$, where $A_{1}$ is a two-qubit register and $A_{2}$ is a single-qubit register.
Now we proceed to calculate the coherent information quantities under the partial decode-forward bound given in Theorem 3. For the coherent information to Bob, we have by the data processing inequality for the coherent information (see Theorem 11.9.3 in Ref. [63]):
$I(A_{1}\rangle B^{\prime}B^{\prime\prime})_{\omega}\geq I(A_{1}\rangle B^{\prime})=(1-2\alpha)\log_{2}4,$ (18)
and for the coherent information to the relay,
$I(A_{1}\rangle E)_{\omega}=(1-2q)\log_{2}4,$ (19)
with $q=\alpha+(1-\alpha)\beta$. Finally, the “conditional term" is bounded as
$I(A_{2}\rangle B^{\prime}B^{\prime\prime}A_{1})\geq I(A_{2}\rangle B^{\prime\prime})=1-2\gamma.$ (20)
Considering the partial decode-forward rate in Eq. (12), we obtain:
$\mathsf{Q}_{\text{PD-F}}(\mathcal{N})\geq\min\bigl{\{}(1-2\alpha)\log_{2}4\,,\\ \;(1-2q)\log_{2}4\bigr{\}}+(1-2\gamma).$ (21)
We note that $q\geq\alpha$, hence the first term reduces to 2(12q)2(1-2q).
Based on Theorem 3, the entanglement-generation capacity of the quantum relay channel above thus satisfies
$C_{\text{EG}}(\mathcal{N})\geq 2(1-2q)+(1-2\gamma),$ (22)
and we can see that we benefit from dividing the transmission between the repeater-aided and the repeaterless links.
VII Summary and Discussion
We study quantum information transmission over a three-terminal quantum relay channel, focusing on decode-forward coding strategies, in both entanglement-assisted and unassisted settings, with either full or partial decoding at the relay. Our model captures key tradeoffs between repeater-aided and repeaterless communications, and highlights the role of entanglement between the transmitter and the relay. Furthermore, our framework generalizes previous results, which focus on either repeaterless transmission or repeater-aided communication separately [28, 29, 67, 51, 42]. These insights serve as a stepping stone towards the design of quantum networks (in the long-term, a quantum Internet), and provide a foundation for practical quantum repeaters and long-distance QKD.
One avenue for future research is to develop alternative coding strategies. In particular, it would be interesting to explore schemes similar to those considered in Ref. [42], such as measure-forward, assist-forward, or compress-and-forward  [68], extending beyond the partial and full decode-forward strategies. Investigating these approaches for quantum channels with quantum messages could reveal new trade-offs between relay-assisted and direct transmission, potentially improving achievable rates or reducing entanglement consumption. Another promising direction is to apply our results to bosonic channels  [69, 70], which model optical communication systems. Extending partial and full decode-forward strategies to bosonic modes could provide insight into achievable rates in realistic continuous-variable quantum networks, including fiber-optic and free-space links.
On a more practical side, when considering how quantum relay channels can be used in the context of distributed relays that facilitate high-rate quantum error correction, it is interesting to analyze our results for quantum channels widely used to model noise in quantum hardware [71, 72]. For example, even local Markovian noise models, such as non-uniform depolarization, generalized amplitude damping, and coherent two-qubit error models often arising from controlled-X gates used in variational quantum algorithms, may provide interesting insights [71, 72, 73]. Analyzing these channels would bridge our primarily quantum communication-focused results with the practical realm of quantum computation.